A jet plane lands with a speed of 100m/s and can accelerate uniformly at a maximum rate of -0.5 m/s2 as it comes to a rest. Can this plane land at an airport where the airport runway is 800 m long. ( hint find the distance it will take to come to rest and compare that to the length of the runway)
1
Expert's answer
2012-09-27T12:17:02-0400
The time needed the plane to stop is

T = V/|a| = 100[m/s]/(|-0.5[m/s²]|) = 200 s,

therefore the stop distance is

D = |a|T²/2 = 0.5[m/s²]*(200[s])²/2 = 10000 m > 800 m.

So, a plane can't land at an airport.